# Light Curve Classification with DistClassiPy: a new distance-based classifier

### v1

@@ Appendix A Distance: A Mathematical Perspective
From the three axioms in Definition 2.1, another condition can be shown to hold:
@@
The 2-dimensional $\mathbb{R}^{2}$ plane with the Euclidean distance is an example of a metric space. Similarly, we can also define metrics on matrices, functions, sets of points or any other mathematical object as long as it satisfies the Definition A.1.
@@ Appendix B List of Metrics
We used the following metrics for all distance calculations in this paper. For a detailed description of each metric, see (Deza and Deza, 2013).
@@
    $\frac{\sum\left(u_{i}-\bar{u})\cdot(v_{i}-\bar{v})\right)}{\sqrt{\sum\left(u_{i}-\bar{u})\cdot(u_{i}-\bar{u})\right)}\sqrt{\sum\left(v_{i}-\bar{v})\cdot(v_{i}-\bar{v})\right)}}$
@@
    Motyka metric:
    The Motyka metric has a summation term in the denominator, which acts as a scaling factor of the maximum term in the numerator.
  13. 13.
@@
    The Correlation metric is based on the Pearson correlation coefficient between 2 points, it is given by:
  18. 18.
@@
    The Maryland Bridge metric is based on the Maryland Bridge similarity, typically used in genomics to measure similarity between genomes.
@@ Appendix C Hardware and Software specifications
The results described in this work, including the computational time estimates, were produced on machines with the hardware specifications reported in Table 4.
@@
The Python libraries and corresponding versions used in this work are :
  • 1.
    Python (Van Rossum and Drake, 2009) version 3.11.8
  • 2.
    DistClassiPy version 0.1.0
  • 3.
    JSON (Pezoa et al., 2016) version 0.13.0
  • 4.
    matplotlib (Hunter, 2007) version 3.8.2
  • 5.
    NumPy (Harris et al., 2020) version 1.26.4
  • 6.
    Pandas (Wes McKinney, 2010; pandas development team, 2020) version 2.2.0
  • 7.
    scikit-learn (Pedregosa et al., 2011) version 1.4.1.post1
  • 8.
    SciPy (Virtanen et al., 2020) version 1.12.0
  • 9.
    seaborn (Waskom, 2021) version 0.13.0
  • 10.
    mlxtend (Raschka, 2018) version 0.23.1
Additional software include Jupyter (Kluyver et al., 2016), tqdm, statistical-distances (Zielezinski, 2023)

### v2

@@ Appendix A Distance: A Mathematical Perspective
From the three axioms in Definition 2.1, another condition (non-negativity) can be shown to hold:
@@
The 2-dimensional $\mathbb{R}^{2}$ plane with the Euclidean distance is an example of a metric space. Similarly, we can also define metrics on matrices, functions, sets of points or any other mathematical object as long as it satisfies A.1.
@@ Appendix B List of Metrics
We used the following metrics for all distance calculations in this paper. For a detailed description of each metric, see Deza and Deza (2013).
@@
    $\frac{\sum\left((u_{i}-\bar{u})\cdot(v_{i}-\bar{v})\right)}{\sqrt{\sum\left((u_{i}-\bar{u})\cdot(u_{i}-\bar{u})\right)}\sqrt{\sum\left((v_{i}-\bar{v})\cdot(v_{i}-\bar{v})\right)}}$
@@
    The Correlation metric is based on the Pearson correlation coefficient between two vectors.
  17. 17.
@@
    The Maryland Bridge metric is based on the Maryland Bridge similarity. This metric does not satisfy axiom 3 (triangle inequality) in Definition 2.1 and it is therefore not a distance metric. However, it has been used in literature, particularly in the field of genomics to measure similarity between genomes (Mirkin and Koonin, 2003). Thus, it is included as a utility in our package.
  18. 18.
    Motyka metric:
    The Motyka metric has a summation term in the denominator, which acts as a scaling factor of the maximum term in the numerator. This metric does not satisfy axiom 1 (identity of indiscernibles) in Definition 2.1 (and it is not positive defined). However, it is a metric that commonly appears in the literature when measuring similarity between data, so it is included as a utility in our package.
@@ Appendix C Hardware and Software specifications
The results described in this work, including the computational time estimates, were produced on machines with the hardware specifications reported in Table 4. The Python libraries and corresponding versions used in this work are given in Table 5.
Specification Details
@@
Library Version Reference
Python 3.12.2 Van Rossum and Drake (2009)
DistClassiPy 0.1.5 Chaini et al. (2024)
NumPy 1.26.4 Harris et al. (2020)
Pandas 2.2.1 Wes McKinney (2010); pandas development team (2020)
scikit-learn 1.4.1.post1 Pedregosa et al. (2011)
SciPy 1.12.0 Virtanen et al. (2020)
statistical-distances 0.9.1 Zielezinski (2023)
matplotlib 3.8.3 Hunter (2007)
JSON 2.0.9 Pezoa et al. (2016)
tqdm 4.65.0 Casper da Costa-Luis et al. (2024)
seaborn 0.13.2 Waskom (2021)
mlxtend 0.23.1 Raschka (2018)
Jupyter 4.1.2 Kluyver et al. (2016)
Table 5: Software and Versions